A neutron star is the remnant left after certain supernovae (explosions of giant stars). Typically, neutron stars are about 16 km in diameter and have around the same mass as our sun. What is a typical neutron star density in g/cm3?

Express your answer using two significant figures.
Answer #1

Mass of our Sun = Mass of the neutron star = M = 1.99 x 1030 kg = 1.99 x 1033 g

Diameter of the neutron star = D = 16 km = 16 x 105 cm

Radius of the neutron star = R = D/2 = 8 x 105 cm

Volume of the neutron star = V

47 R3 V =

Density of the neutron star = \rho

\rho = \frac{M}{V}

V P (47Rº/3)

ЗМ pe AR

3(1.99 + 103) -4718 + 10553

\rho = 9.3 x 1014 g/cm3

Typical neutrons star density = 9.3 x 1014 g/cm3